The smallest angle in a triangle is 1/2 as large as the largest angle
The third angle is twice the smallest angle. Find the three angles.

Answer by checkley79(3341)

SMALLEST ANGLE=L/2
THE THIRD ANGLE=2S
S+L+T=180
L/2+L+2(L/2)=180
L/2+L+L=180
2.5L=180
L=180/2.5
L=72 DEGREES FOR THE LARGEST ANGLE.
S=72/2=36 DEGREES FOR THE SMALLEST ANGLE.
2*36=72 DEGREES FOR THE THIRD ANGLE.
PROOF:
72+36+72=180
180=180